OpenStudy (anonymous):

simplify: (x2-4)/(2-x)

OpenStudy (anonymous):

need help again

OpenStudy (anonymous):

[ (x+2)(x-2) ] / (2-x) the top is the same as -[ (-2-x)(2-x) ] so the bottom 2-x cancels out and you're left with -(-2-x) which is just 2+x

myininaya (myininaya):

(x+2)(x-2)=x^2-4 {(x+2)(x-2)}/(2-x) but 2-x=-1(x-2) so we have {(x+2)(x-2)}/(-1(x-2)) x-2 cancels (x+2)/(-1) -1(x+2) -(x+2)

OpenStudy (anonymous):

so the answer is -(x+2) ?

myininaya (myininaya):

yes the answer is -x-2 or -(x+2)

myininaya (myininaya):

ivo you can canceled -2-x with 2-x these are different

myininaya (myininaya):

cant't *

OpenStudy (anonymous):

oh, i see.. k good to know. thanks.